# RLTHF: Targeted Human Feedback for LLM Alignment

### v2

@@ Abstract
Fine-tuning large language models (LLMs) to align with user preferences is challenging due to the high cost of quality human annotations in Reinforcement Learning from Human Feedback (RLHF) and the generalizability limitations of AI Feedback. To address these challenges, we propose RLTHF, a human-AI hybrid framework that combines LLM-based initial alignment with selective human annotations to achieve full-human annotation alignment with minimal effort. RLTHF identifies hard-to-annotate samples mislabeled by LLMs using a reward model’s reward distribution and iteratively enhances alignment by integrating strategic human corrections while leveraging LLM’s correctly labeled samples. Evaluations on HH-RLHF and TL;DR datasets show that RLTHF reaches full-human annotation-level alignment with only 6-7% of the human annotation effort. Furthermore, models trained on RLTHF’s curated datasets for downstream tasks outperform those trained on fully human-annotated datasets, underscoring the effectiveness of RLTHF’s strategic data curation.
[…]
*Work is done during an internship at Microsoft Research.
@@ 1 Introduction
Figure 1: Overview of RLTHF process. RLTHF starts with coarse LLM alignment on the task. It then iteratively takes targeted human feedback to reach the complete human alignment, leveraging reward distribution of a reward model in its training dataset.
@@
First, we develop a concept that leverages the reward distribution of a reward model over its training dataset to capture the relative arrangement of samples based on rewarded features. This property allows us to detect plausible inaccuracies in annotations across the dataset. Specifically, we train a reward model on the LLM-labeled dataset to identify clusters of hard-to-annotate samples that are highly likely to be either mislabeled or correctly labeled by the LLM.
Building on this concept, we propose an innovative iterative reward model training technique to achieve oracle-level human alignment in the dataset. In each iteration, RLTHF identifies and rectifies highly probable mislabeled data points using human annotations. Simultaneously, it detects clusters of samples that are very likely to be correctly labeled by the LLM and incorporates them with human-annotated data to construct a high-quality training set for the next iteration of reward model training. Throughout this process, RLTHF preserves data richness and maximizes the efficiency of human annotation investment through carefully controlled hyperparameters.
Finally, we implement and evaluate RLTHF on two distinct preference datasets: HH-RLHF and TL;DR. Our results demonstrate that RLTHF achieves accuracy comparable to a fully human-annotated dataset while requiring only 6–7% of the total human annotations. Furthermore, we conduct a comparative study by training models on downstream tasks using DPO. Remarkably, models trained with RLTHF even outperform those trained on fully human-annotated datasets, highlighting the impact of RLTHF’s meticulous data curation in enhancing model performance.
@@ 2.2 LLM Self-Improvement
To break the upper bound of LLMs, recent efforts have been devoted to enabling LLMs to self-improve. Self-Rewarding LMs (Yuan et al., 2024) and Math-shepherd (Wang et al., 2024b) demonstrate the possibility of LLM self-improvement with reward signals from itself. SELF-ALIGN (Sun et al., 2024b) uses a carefully written set of principles to guide LLMs through self-improvement. SER (Huang et al., 2024) starts with only a fraction of human annotations to achieve full-annotation performance by progressively generating additional training data for itself. However, these methods still suffer from the intrinsic upper bound of LLMs and self-improvement is not guaranteed for customized tasks. RLTHF, on the other hand, efficiently introduces new human intelligence into the improvement process, thereby ensuring that the improvement is not bounded by LLMs’ initial lack of domain understanding.
@@ 3 Improving Human Alignment with RLTHF
RLTHF enhances human alignment in preference datasets used for training preference optimization techniques like DPO and PPO. It facilitates LLM training for various downstream tasks, including summarization, compliance, and grounding. Starting with an unlabeled preference dataset, RLTHF strategically integrates AI-generated labels with selective human feedback to maximize alignment while minimizing annotation effort. As illustrated in Figure 1, RLTHF operates in three stages: 1) Initial alignment, where an off-the-shelf LLM provides dataset labeling to establish a coarse task understanding, 2) Iterative alignment improvement, which leverages reward distribution by an RM to locate and rectify hard-to-annotate samples mislabeled by the LLM with selective human feedback while investing the correct LLM labels, 3) Transferring knowledge for downstream task, where the curated preference dataset is fed into the DPO pipeline or the trained RLTHF reward model is integrated into the PPO pipeline.
@@ 3.1 Initial Alignment
Our prompt for obtaining preference judgments from the LLM consists of three key components: 1) task description, 2) preference judgment principles provided by the end user, and 3) few-shot examples with optional chain-of-thought (CoT) reasoning. The prompt templates are detailed in Appendix A. We do not perform explicit fine-grained prompt tuning, as full visibility into the data may be restricted when offering fine-tuning services to third-party customers. However, to ensure that the selected LLM with our prompt attains a reasonable level of alignment, we perform an eyes-off validation using strategic human feedback, as detailed in Section 3.2.3.
As mentioned earlier, AI-generated feedback is prone to errors due to factors such as model biases from pre-training data, task complexity, and prompt optimization, which is also evident in our evaluation. When our ultimate goal is to customize an existing model through fine-tuning to align with end-user preferences, we inherently assume that an off-the-shelf LLM lacks comprehensive alignment with the end-user. However, RLTHF builds upon the initial AI-provided alignment and systematically refines it in subsequent stages to achieve oracle-level human alignment.
@@ 3.2 Iterative Alignment Improvement
In this stage, we refine the LLM-labeled preference dataset by iteratively training a reward model (RM) with selective human annotations to enhance alignment. Before diving into the details of this process, we first establish the premise for RM.
@@ 3.2.1 Reward Model
Given a labeled preference dataset $\mathcal{D}_{\mathbf{\Lambda}}=\{x_{i},y_{i,c},y_{i,r}\}$, where $i\in[N]$, $x_{i}$ is the prompt, $y_{i,c}$ and $y_{i,r}$ denote the chosen and rejected completions, respectively, as labeled according to the annotator’s preference, $\mathbf{\Lambda}$. Here, if we represent the relative preference orientation of $i^{th}$ completion pair with $\lambda=[-1,+1]$, $\mathbf{\Lambda}$ is a $N$-dimensional vector consists of $[\lambda_{i}]_{i=1}^{N}$, meaning that flipping the preferences of all completion pairs results in $\mathcal{D}_{\mathbf{-\Lambda}}$. To train an RM on this dataset, we can formulate the probability distribution of $y_{i,c}$ being preferred over $y_{i,r}$ given $x_{i}$ as an input, following the Bradley-Terry (BT) model (David, 1963).
@@ 3.2.2 Looking at Reward Distribution
The reward distribution curve $\vartheta(\Delta_{\mathbf{\Lambda}}{\hat{r}_{\mathbf{\Lambda}})}$, derived from the training preference dataset $\mathcal{D}_{\mathbf{\Lambda}}$, reflects the degree of alignment the RM (trained with optimal validation loss) has achieved during training across $\mathcal{D}_{\mathbf{\Lambda}}$. The upper left region of the curve consists of samples with high positive $\Delta_{\mathbf{\Lambda}}\hat{r}_{\mathbf{\Lambda}}$, indicating strong agreement between the RM and the training preference labels $\mathbf{\Lambda}$. This suggests that the RM effectively identifies and reinforces strong winning preference features in these samples, implying that these features were dominant in $\mathcal{D}_{\mathbf{\Lambda}}$. Conversely, the bottom right region of the curve contains samples with very low or even negative $\Delta_{\mathbf{\Lambda}}\hat{r}_{\mathbf{\Lambda}}$, signaling disagreement between the trained reward function $\hat{r}_{\mathbf{\Lambda}}$ and the training preference label for these samples. This misalignment arises from two primary factors. (1) Absence of strong features, where RM is not able to find any strong preference feature in these samples according to $\hat{r}_{\mathbf{\Lambda}}$. (2) Conflicting samples within $\mathcal{D}_{\mathbf{\Lambda}}$, where the preference features of these samples are highly conflicting with other stronger preference features learned in $\hat{r}_{\mathbf{\Lambda}}$, leading the RM to penalize them.
@@ 3.2.3 RLTHF Leveraging Reward Distribution
RLTHF trains the initial RM using a preference dataset filtered by a general-purpose LLM in the previous stage. We denote this dataset as $\mathcal{D}_{\mathbf{\Lambda_{LLM}}}$ where $\mathbf{\Lambda}_{LLM}$ represents the preference labeling performed by the LLM. Since the RM training includes a validation set derived from $\mathcal{D}_{\mathbf{\Lambda_{LLM}}}$, this ensures that the trained RM is broadly aligned with the LLM’s preferences. We assume that the LLM has a coarse but reasonable understanding of preference judgments, particularly for relatively easy-to-annotate samples. As a result, the features of these samples dominate in $\hat{r}_{\mathbf{\Lambda_{LLM}}}$. Based on our earlier discussion, the upper left region of the reward density curve, $\vartheta(\Delta_{\mathbf{\Lambda_{LLM}}}{\hat{r}_{\mathbf{\Lambda_{LLM}}})}$ contains high density of samples with prominent preference features, i.e., those that are easier for the LLM to annotate accurately. Before proceeding, we further validate that the LLM is at least roughly aligned with the end-user’s preferences in terms of these easy-to-annotate samples. This step mitigates the risk of significant misalignment due to prompt curation or model selection. To achieve this, RLTHF automatically (details in the following section) samples a small subset ($<0.1\%$) of preference data from the upper left region and gathers eyes-off user feedback. If human agreement on these samples is low, it signals a major misalignment between the user’s preferences and the LLM’s outputs. While we did not observe such cases in our experiments, this issue can be addressed by refining the judgment principles in the prompt. Updates can be made directly by the user, by incorporating verbose user feedback, or even through automated prompt optimization techniques (Kepel & Valogianni, 2024; Li et al., 2024).
@@
At this stage, we can identify regions with a high density of correctly labeled samples by the LLM, i.e., those that are relatively easy for the LLM to annotate in alignment with human preference. Now, we turn our attention to two critical types of samples necessary for achieving fine-grained alignment: (1) hard-to-annotate samples and (2) samples mislabeled by the LLM w.r.t. the human preference $\mathbf{\Lambda}_{h}$. Since the LLM was unable to correctly label these samples initially, the reward function $\hat{r}_{\mathbf{\Lambda_{LLM}}}$ cannot accurately capture their preference features. Consequently, these samples are expected to cluster around the bottom right region of the reward distribution curve $\vartheta(\Delta_{\mathbf{\Lambda_{LLM}}}{\hat{r}_{\mathbf{\Lambda_{LLM}}})}$. To illustrate this, we refer to Figure 2(a) and 2(b). Figure 2(a) shows $\vartheta(\Delta_{\mathbf{\Lambda_{LLM}}}{\hat{r}_{\mathbf{\Lambda_{LLM}}})}$ from one of our experiments. In this figure, we classify each sample $\rho_{i}\in\mathcal{D}_{\mathbf{\Lambda_{LLM}}}$ as either correctly or incorrectly labeled w.r.t. the human preference $\mathbf{\Lambda_{h}}$, i.e., whether the preference assigned by $\mathbf{\Lambda_{LLM}}$ is matching $\mathbf{\Lambda_{h}}$. As observed, the upper left region of the curve contains a high density of correctly labeled samples, supporting our earlier claim that these represent the LLM’s easy-to-annotate cases. To quantify this, we generate an accuracy density curve for $\vartheta(\Delta_{\mathbf{\Lambda_{LLM}}}{\hat{r}_{\mathbf{\Lambda_{LLM}}})}$ w.r.t. the full-human preference $\mathbf{\Lambda_{h}}$, as shown in Figure 2(b). This figure confirms that alignment with human preference decreases as we move towards the right side of the curve.
@@ 3.2.4 Selective Human Annotation
To enhance alignment from this stage, human annotation is necessary, but it must be done efficiently to maximize its impact. A straightforward approach is to refer to the accuracy density curve—annotations in the lowest accuracy region will yield the highest benefit. Thus, we could start annotating from the very bottom of the curve. However, as previously discussed, some samples in this region may exhibit preference features that are largely opposite to the dominant features captured by $\hat{r}_{\mathbf{\Lambda_{LLM}}}$. These samples are highly likely to be mislabeled in $\mathbf{\Lambda_{LLM}}$ (see Appendix B). Instead of requiring human annotation, we can simply flip the preference of these samples to correct the mislabeling. To estimate the positions of such samples, we take the reflection of the “elbow” point across the x-axis, as the elbow marks the region containing strong preference features. This reflected point, known as the “reflection point”, always lies to the right of the “knee” in the lowest accuracy density region. Human annotation then begins from the reflection point, ensuring the most effective correction of alignment errors.
@@ 3.2.5 Iterative Approach
The current reward function $\hat{r}_{\mathbf{\Lambda_{LLM}}}$, trained on $\mathcal{D}_{\mathbf{\Lambda_{LLM}}}$, exhibits an alignment gap with respect to $\mathbf{\Lambda_{h}}$ due to the presence of hard-to-annotate samples for the LLM and mislabeling by the LLM. Since we have identified ways to rectify these issues, we can refine $\mathcal{D}$ to improve alignment and train a new RM that better aligns with $\mathbf{\Lambda_{h}}$. Now, the question is how to prepare the dataset for the next iteration of RM training. Suppose we are currently in iteration 0 (Itr-0) with $\mathcal{D}_{\mathbf{\Lambda_{LLM}}}$ and $\hat{r}_{\mathbf{\Lambda_{LLM}}}$. For the iteration 1 (Itr-1) training dataset, $\mathcal{D}_{\mathbf{\Lambda_{Itr-1}^{T}}}^{T}$, our primary goal is to include high-confidence samples that are well-aligned with $\mathbf{\Lambda_{h}}$. The first choice is definitely human annotated samples from Itr-0. Additionally, another set of candidates can be drawn from the high-accuracy density region of $\vartheta(\Delta_{\mathbf{\Lambda_{LLM}}}{\hat{r}_{\mathbf{\Lambda_{LLM}}})}$, specifically the region to the left of the “elbow”, where the RM has learned strong preference features in alignment with $\mathbf{\Lambda_{h}}$.
Although these two sets of samples offer high precision, $\mathcal{D}_{\mathbf{\Lambda_{Itr-1}^{T}}}^{T}$ will still face a data coverage issue. These two candidate sets represent samples with the longest feature distance, leaving gaps in intermediate regions. However, expanding the dataset by including samples from the middle region, i.e., right of the “elbow” and left of the “knee” risks introducing misaligned samples. Since the accuracy in this region is likely to be just above $50\%$, obtaining human annotations for these samples would be inefficient. Furthermore, as the number of samples annotated from the right of the knee is relatively small, their preference features are likely to be overshadowed by the dominant preference features of the high numbers of left-side samples. As a result, their features may not be effectively captured in $\hat{r}_{\mathbf{\Lambda_{Itr-1}^{T}}}$. To balance these trade-offs, we introduce two hyperparameters specific to RLTHF, allowing for a more controlled and effective dataset expansion while maintaining alignment quality.
@@
The dataset $\mathcal{D}_{\mathbf{\Lambda_{Itr-1}}}^{T}$ consists of carefully selected samples from $\mathcal{D}_{\mathbf{\Lambda_{LLM}}}$, ensuring high alignment with $\mathbf{\Lambda_{h}}$ by optimally tuning the hyperparameters $\alpha$ and $\beta$. Training the RM on $\mathcal{D}_{\mathbf{\Lambda_{Itr-1}^{T}}}^{T}$ results in $\hat{r}_{\mathbf{\Lambda_{Itr-1}^{T}}}$, which is more closely aligned with $\mathbf{\Lambda_{h}}$. After training, we construct the dataset for generating the reward distribution curve by incorporating the remaining samples from Itr-0: $\mathcal{D}_{\mathbf{\Lambda_{Itr-1}}}=\mathcal{D}_{\mathbf{\Lambda_{Itr-1}^{T}}}^{T}\cup(\mathcal{D}_{\mathbf{\Lambda_{LLM}}}-\mathcal{D}_{\mathbf{\Lambda_{Itr-1}^{T}}}^{T})$. From this, we generate a new reward distribution curve, $\vartheta(\Delta_{\mathbf{\Lambda_{Itr-1}}}{\hat{r}_{\mathbf{\Lambda_{Itr-1}^{T}}})}$. While this curve demonstrates improved alignment with $\mathbf{\Lambda_{h}}$, full alignment is not necessarily achieved. However, it presents RLTHF with a distinct reward distribution curve compared to the previous iteration. This evolving diversity in $\vartheta(\cdot)$ enhances the variety of human annotations, maximizing the return on annotation investments and incrementally enriching $\mathcal{D}$. It is important to note that the effectiveness of this diversification, as well as the corresponding improvements, depends on factors such as hyperparameter tuning (see Section 4), the original data distribution, and model selection.
As we have seen, RLTHF maximizes the efficiency of human annotations by iteratively refining $\vartheta(\cdot)$ and exposing annotators to diverse, LLM-mislabeled samples. To further enhance annotation efficiency, RLTHF employs random sharding to down-sample the original corpus. It begins by selecting a random shard of the dataset, iteratively improving alignment within that subset. Once the desired alignment is achieved, the final iteration’s RM is used to label the entire corpus. This approach enables RLTHF to concentrate human annotations in a smaller, more targeted space while effectively propagating alignment across the full dataset at the end.

### v3

@@ Abstract
Fine-tuning large language models (LLMs) to align with user preferences is challenging due to the high cost of quality human annotations in Reinforcement Learning from Human Feedback (RLHF) and the generalizability limitations of AI Feedback. To address these challenges, we propose RLTHF, a human-AI hybrid framework that combines LLM-based initial alignment with selective human annotations to achieve full-human annotation alignment with minimal effort. RLTHF identifies hard-to-annotate samples mislabeled by LLMs using a reward model’s reward distribution and iteratively enhances alignment by integrating strategic human corrections while leveraging LLM’s correctly labeled samples. Evaluations on HH-RLHF and TL;DR datasets show that RLTHF reaches full-human annotation-level alignment with only 6-7% of the human annotation effort. Furthermore, models trained on RLTHF’s curated datasets for downstream tasks outperform those trained on fully human-annotated datasets, underscoring the effectiveness of RLTHF.
LLM, RLHF, Alignment, Machine Learning, ICML
@@ 1 Introduction
Figure 1: Overview of RLTHF process. RLTHF starts with coarse LLM alignment on the task. It then iteratively takes targeted human feedback and blending the same with sanitized LLM-annotations to reach the complete human alignment, leveraging reward distribution of a reward model in its training dataset.
@@
First, we introduce a concept that leverages the reward distribution of a reward model over its training dataset to capture the relative arrangement of samples based on rewarded features. This distributional property enables the identification of both potential annotation inaccuracies and the model’s confidence across the dataset. Specifically, we train a reward model on the LLM-labeled dataset to uncover clusters of hard-to-annotate samples that are highly likely to be mislabeled by the LLM, as well as easy samples that are highly likely to be correctly labeled.
Building on this concept, we propose an innovative iterative reward model training technique to achieve oracle-level human alignment in the dataset. In each iteration, RLTHF identifies highly probable mislabeled data points and rectifies the same using human annotations. Simultaneously, it detects clusters of samples that are very likely to be correctly labeled by the LLM and incorporates them with human-annotated data to construct a high-quality training set for the next iteration of reward model training. Throughout this process, RLTHF preserves data richness and maximizes the efficiency of human annotation investment through carefully controlled hyperparameters.
Finally, we evaluate RLTHF on two distinct preference datasets: HH-RLHF and TL;DR. Our results demonstrate that RLTHF achieves accuracy comparable to a fully human-annotated dataset while requiring only 6–7% of the total human annotations. Furthermore, we conduct a comparative study by training models on downstream tasks using DPO (Rafailov et al., 2024). Remarkably, models trained with RLTHF even outperform those trained on fully human-annotated datasets, highlighting the impact of RLTHF’s meticulous data curation in enhancing model performance.
@@ 2.2 LLM Self-Improvement
To break the upper bound of LLMs, recent efforts have been devoted to enabling LLMs to self-improve. Self-Rewarding LMs (Yuan et al., 2024) and Math-shepherd (Wang et al., 2024b) demonstrate the possibility of LLM self-improvement with reward signals from itself. SELF-ALIGN (Sun et al., 2024b) uses a carefully written set of principles to guide LLMs through self-improvement. SER (Huang et al., 2024) starts with only a fraction of human annotations to achieve full-annotation performance by progressively generating additional training data for itself. However, these methods still suffer from the intrinsic upper bound of LLMs and self-improvement is not guaranteed for customized tasks. RLTHF, on the other hand, efficiently introduces human intelligence into the improvement process, thereby ensuring that the improvement is not bounded by LLMs’ initial lack of domain understanding.
@@ 3 Improving Human Alignment with RLTHF
RLTHF enhances alignment with human in preference datasets used for training preference optimization techniques like DPO and PPO. It facilitates LLM training for various downstream tasks, including summarization, compliance, and grounding. Starting with an unlabeled preference dataset, RLTHF strategically integrates AI-generated labels with selective human feedback to maximize alignment while minimizing annotation effort. As illustrated in Figure 1, RLTHF operates in three stages: 1) Initial alignment, where an off-the-shelf LLM provides dataset labeling to establish a coarse task understanding, 2) Iterative alignment improvement, which leverages reward distribution by a reward model (RM) to locate hard-to-annotate samples mislabeled by the LLM and rectify with selective human feedback while investing the correct LLM labels, 3) Transferring knowledge for downstream task, where the curated preference dataset is fed into the DPO pipeline or the trained RLTHF reward model is integrated into the PPO pipeline. Find the corresponding pseudocode in Appendix B.
@@ 3.1 Initial Alignment
Our prompt for obtaining preference judgments from the LLM consists of three components: 1) task description, 2) preference judgment principles provided by the user, and 3) few-shot examples with optional chain-of-thought reasoning. The prompt templates are detailed in Appendix A. We do not perform explicit fine-grained prompt tuning, as full visibility into the data may be restricted when offering fine-tuning services to third-party customers. However, to ensure that the selected LLM with our prompt attains a rough level of alignment, we perform an eyes-off validation using strategic human feedback, as detailed in Section 3.2.3.
As mentioned earlier, this AI-generated feedback is prone to errors due to factors such as model biases from pre-training data, task complexity, and prompt optimization, which is also evident in our evaluation. When our ultimate goal is to customize an existing model through fine-tuning to align with end-user preferences, we inherently assume that an off-the-shelf LLM lacks comprehensive alignment with the end-user. However, RLTHF builds upon the initial AI-provided alignment and systematically refines it in subsequent stages to achieve oracle-level human alignment.
@@ 3.2 Iterative Alignment Improvement
In this stage, we refine the LLM-labeled preference dataset by iteratively training an RM with selective human annotations to enhance alignment. Before diving into the details of this process, we first establish the premise for RM.
@@ 3.2.1 Reward Model
Given a labeled preference dataset $\mathcal{D}_{\mathbf{\Lambda}}=\{x_{i},y_{i,c},y_{i,r}\}$, where $i\in[N]$, $x_{i}$ is the prompt, $y_{i,c}$ and $y_{i,r}$ denote the chosen and rejected completions, respectively, as labeled according to the annotator’s preference, $\mathbf{\Lambda}$. Here, if we represent the relative preference orientation of $i^{th}$ completion pair with $\lambda=[-1,+1]$, $\mathbf{\Lambda}$ is a $N$-dimensional vector consists of $\lambda_{i}]_{i=1}^{N}$, meaning that flipping the preferences of all completion pairs results in $\mathcal{D}_{\mathbf{-\Lambda}}$. To train an RM on this dataset, we can formulate the probability distribution of $y_{i,c}$ being preferred over $y_{i,r}$ given $x_{i}$ as an input, following the Bradley-Terry (BT) model (David, 1963).
@@ 3.2.2 Looking at Reward Distribution
The reward distribution curve $\vartheta(\Delta_{\mathbf{\Lambda}}{\hat{r}_{\mathbf{\Lambda}})}$, derived from the training preference dataset $\mathcal{D}_{\mathbf{\Lambda}}$, reflects the degree of alignment the RM (trained with optimal validation loss) has achieved across $\mathcal{D}_{\mathbf{\Lambda}}$ during training. The upper left region of the curve consists of samples with high positive $\Delta_{\mathbf{\Lambda}}\hat{r}_{\mathbf{\Lambda}}$, indicating strong agreement between the RM and the training preference labels $\mathbf{\Lambda}$. This suggests that the RM effectively identifies and reinforces strong winning preference features in these samples, implying that these features were dominant in $\mathcal{D}_{\mathbf{\Lambda}}$. Conversely, the bottom right region of the curve contains samples with very low or even negative $\Delta_{\mathbf{\Lambda}}\hat{r}_{\mathbf{\Lambda}}$, signaling disagreement between the trained reward function $\hat{r}_{\mathbf{\Lambda}}$ and the training preference labels for these samples. This misalignment arises from two primary factors. (1) Absence of strong features, where RM is not able to find any strong preference feature in these samples according to $\hat{r}_{\mathbf{\Lambda}}$. (2) Conflicting samples within $\mathcal{D}_{\mathbf{\Lambda}}$, where the preference features of these samples are highly conflicting with other stronger preference features learned in $\hat{r}_{\mathbf{\Lambda}}$, leading the RM to penalize them.
@@ 3.2.3 RLTHF Leveraging Reward Distribution
RLTHF trains the initial RM using a preference dataset labeled by a general-purpose LLM from the previous stage. We denote this dataset as $\mathcal{D}_{\mathbf{\Lambda_{LLM}}}$ where $\mathbf{\Lambda}_{LLM}$ represents the LLM’s preference labeling. Since the RM training includes a validation set derived from $\mathcal{D}_{\mathbf{\Lambda_{LLM}}}$, this ensures that the trained RM is broadly aligned with the LLM’s preferences. We assume that the LLM has a coarse but reasonable understanding of preference judgments, particularly for relatively easy-to-annotate samples. As a result, the features of these samples dominate in $\hat{r}_{\mathbf{\Lambda_{LLM}}}$. Based on our earlier discussion, the upper left region of the reward density curve, $\vartheta(\Delta_{\mathbf{\Lambda_{LLM}}}{\hat{r}_{\mathbf{\Lambda_{LLM}}})}$ contains high density of samples with prominent preference features, i.e., those that are easier for the LLM to annotate accurately. Before proceeding, we further validate that the LLM is at least roughly aligned with the user in terms of these easy-to-annotate samples. This step mitigates the risk of significant misalignment due to prompt curation or model selection. To achieve this, RLTHF automatically (details in the following section) samples a small subset ($0.1\%$) of preference data from the upper left region and gathers user feedback. If human agreement on these samples is low, it signals a major misalignment between the user and LLM. While we did not observe such cases in our experiments, this issue can be addressed by refining the judgment principles in the prompt. Updates can be directly made by the user, through incorporating verbose user feedback, or even through automated prompt optimization techniques (Kepel & Valogianni, 2024).
@@
At this stage, we can identify regions with a high density of correctly labeled samples by the LLM, i.e., those that are relatively easy for the LLM to annotate in alignment with human preference. Now, we turn our attention to two critical types of samples necessary for achieving fine-grained alignment: (1) hard-to-annotate samples and (2) samples mislabeled by the LLM w.r.t. the human preference $\mathbf{\Lambda}_{h}$. Since the LLM was unable to correctly label these samples initially, the reward function $\hat{r}_{\mathbf{\Lambda_{LLM}}}$ cannot accurately capture their preference features. Consequently, these samples are expected to cluster around the bottom right region of the reward distribution curve $\vartheta(\Delta_{\mathbf{\Lambda_{LLM}}}{\hat{r}_{\mathbf{\Lambda_{LLM}}})}$. To illustrate this, we refer to Figure 2(a) and 2(b). Figure 2(a) shows $\vartheta(\Delta_{\mathbf{\Lambda_{LLM}}}{\hat{r}_{\mathbf{\Lambda_{LLM}}})}$ from one of our experiments. In this figure, we classify each sample $\rho_{i}\in\mathcal{D}_{\mathbf{\Lambda_{LLM}}}$ as either correctly or incorrectly labeled w.r.t. the human preference $\mathbf{\Lambda_{h}}$, i.e., whether the preference assigned by $\mathbf{\Lambda_{LLM}}$ is matching $\mathbf{\Lambda_{h}}$. As observed, the upper left region of the curve contains a high density of correctly labeled samples, supporting our earlier claim that these represent the LLM’s easy-to-annotate cases. To quantify this, we generate an accuracy density curve for $\vartheta(\Delta_{\mathbf{\Lambda_{LLM}}}{\hat{r}_{\mathbf{\Lambda_{LLM}}})}$ w.r.t. the human preference $\mathbf{\Lambda_{h}}$, as shown in Figure 2(b). This figure confirms that alignment with human preference decreases as we move towards the right side of the curve.
@@ 3.2.4 Selective Human Annotation
To enhance alignment from this stage, human annotation is necessary, but it must be done efficiently to maximize its impact. A straightforward approach is referring to the accuracy density curve – annotations in the lowest accuracy region would yield the highest benefit. Thus, we could start annotating from the very bottom of the curve. However, as previously discussed, some samples in this region may exhibit preference features that are largely opposite to the dominant features captured by $\hat{r}_{\mathbf{\Lambda_{LLM}}}$. These samples are highly likely to be mislabeled in $\mathbf{\Lambda_{LLM}}$ (see Appendix C). Instead of seeking human annotation, we can simply flip the preference of these samples to rectify. To estimate the location of such samples, we take the reflection of the “elbow” point w.r.t. the x-axis, as the elbow marks the region containing strong preference features. This “reflection point” always lies to the right of the “knee” in the lowest accuracy density region. We begin human annotation at this “reflection point” and proceed leftward along the curve, ensuring the most effective correction of alignment errors.
@@ 3.2.5 Iterative Approach
The current reward function $\hat{r}_{\mathbf{\Lambda_{LLM}}}$, trained on $\mathcal{D}_{\mathbf{\Lambda_{LLM}}}$, exhibits an alignment gap w.r.t. $\mathbf{\Lambda_{h}}$ due to the presence of hard-to-annotate samples for the LLM and mislabeling by the LLM. Since we have identified ways to rectify these issues, we can refine $\mathcal{D}$ to improve alignment and train a new RM that better aligns with $\mathbf{\Lambda_{h}}$. Now, the question is how to prepare the dataset for the next iteration of RM training? Suppose we are currently in iteration 0 (Itr-0) with $\mathcal{D}_{\mathbf{\Lambda_{LLM}}}$ and $\hat{r}_{\mathbf{\Lambda_{LLM}}}$. For the iteration 1 (Itr-1) training dataset, $\mathcal{D}_{\mathbf{\Lambda_{Itr-1}^{T}}}^{T}$, our primary goal is to include high-confidence samples that are well-aligned with $\mathbf{\Lambda_{h}}$. The first choice is definitely human annotated samples from Itr-0. Additionally, another set of candidates can be drawn from the high-accuracy density region of $\vartheta(\Delta_{\mathbf{\Lambda_{LLM}}}{\hat{r}_{\mathbf{\Lambda_{LLM}}})}$, specifically the region to the left of the “elbow”, where the RM has learned strong preference features in alignment with $\mathbf{\Lambda_{h}}$.
Although these two sets of samples offer high precision, $\mathcal{D}_{\mathbf{\Lambda_{Itr-1}^{T}}}^{T}$ will still face a data coverage issue. Looking at the reward distribution curve, these two candidate sets represent samples with the longest distance, leaving gaps in middle region. However, expanding the dataset by including samples from the middle region, i.e., right of the “elbow” and left of the “knee” risks introducing misaligned samples. Since the accuracy in this region is likely to be just above $50\%$, obtaining human annotations for these samples would be inefficient. Furthermore, as the number of samples annotated from the right of the knee is relatively small, their preference features are likely to be overshadowed by the dominant preference features of the high numbers of left-side samples. As a result, their features may not be effectively captured in $\hat{r}_{\mathbf{\Lambda_{Itr-1}^{T}}}$. To balance these trade-offs, we introduce two hyperparameters, allowing for a more controlled and effective dataset expansion while maintaining alignment quality.
@@
The dataset $\mathcal{D}_{\mathbf{\Lambda_{Itr-1}}}^{T}$ consists of carefully selected samples from $\mathcal{D}_{\mathbf{\Lambda_{LLM}}}$, ensuring high alignment with $\mathbf{\Lambda_{h}}$ by optimally tuning the hyperparameters $\alpha$ and $\beta$. Training the RM on $\mathcal{D}_{\mathbf{\Lambda_{Itr-1}^{T}}}^{T}$ results in $\hat{r}_{\mathbf{\Lambda_{Itr-1}^{T}}}$, which is more closely aligned with $\mathbf{\Lambda_{h}}$. After training, we construct the dataset for generating the reward distribution curve by incorporating the remaining samples from Itr-0: $\mathcal{D}_{\mathbf{\Lambda_{Itr-1}}}=\mathcal{D}_{\mathbf{\Lambda_{Itr-1}^{T}}}^{T}\cup(\mathcal{D}_{\mathbf{\Lambda_{LLM}}}-\mathcal{D}_{\mathbf{\Lambda_{Itr-1}^{T}}}^{T})$. From this, we generate a new reward distribution curve, $\vartheta(\Delta_{\mathbf{\Lambda_{Itr-1}}}{\hat{r}_{\mathbf{\Lambda_{Itr-1}^{T}}})}$. While this curve demonstrates improved alignment with $\mathbf{\Lambda_{h}}$, full alignment is not necessarily achieved. However, it presents RLTHF with a distinct reward distribution curve compared to the previous iteration. This evolving diversity in $\vartheta(\cdot)$ enhances the variety of human annotations, maximizing the return on annotation investments and incrementally enriching $\mathcal{D}$. Note that the effectiveness of this diversification, as well as the corresponding improvements, depends on factors such as hyperparameter tuning (see Section 4), the original data distribution, and model selection.
RLTHF maximizes the efficiency of human annotations by iteratively refining $\vartheta(\cdot)$ and exposing annotators to diverse, LLM-mislabeled samples. To further enhance annotation efficiency, RLTHF employs random sharding to down-sample the original corpus. It begins by selecting a random shard of the dataset, iteratively improving alignment within that subset. Once the desired alignment is achieved, the final iteration’s RM is used to label the entire corpus. This approach enables RLTHF to concentrate human annotations in a smaller, more targeted space while effectively propagating alignment across the full dataset at the end.